# One Image is Worth a Thousand Words: A Usability Preservable Text-Image Collaborative Erasing Framework

### v1

@@ 1 Introduction
Recent years have witnessed significant progress in the field of generation (Kingma & Welling, 2013; Goodfellow et al., 2014; Mirza & Osindero, 2014; Sohn et al., 2015), especially the diffusion model (Ho et al., 2020; Song et al., 2021). For example, the popular stable diffusion model  [^1] https://github.com/CompVis/stable-diffusion, trained on the LAION-5B dataset (Schuhmann et al., 2022), can produce high-quality images aligned with a given text prompt and have overwhelmed the area of image reconstruction (Takagi & Nishimoto, 2023; Wu et al., 2024; Li et al., 2023), edition (Meng et al., 2022; Yang et al., 2023; Kawar et al., 2023; Bao et al., 2019) and generation (Rombach et al., 2022; Ho et al., 2022; Zhou et al., 2023). Despite its impressive generation ability, the diffusion model also raises safety concerns, one of which is the potential to generate unwanted content, such as NSFW (Not Safe For Work) images. Taking the SD v1.4 in  LABEL:fig:intro_example as an example, the designed prompt is semantically benign but induces explicit output.
To address this issue, a wide range of studies (Gandikota et al., 2023; Zhang et al., 2024a; Wu & Harandi, 2024; Schramowski et al., 2023; Kumari et al., 2023; Gandikota et al., 2024; Lyu et al., 2024; Zhang et al., 2024b; Fan et al., 2024; Huang et al., 2023; Bui et al., 2024; Changhoon et al., 2024; Chao et al., 2024; Lu et al., 2024; Alvin & Harold, 2023) have explored concept erasing techniques, which rely solely on text as guidance. However, as shown in this paper, their performances are limited by the inherent gap between the text and image modalities. Concepts are often complex and deeply entangled, making it challenging to fully separate with only text descriptions. Therefore, merely using text easily results in an incomplete erasing and affects other untargeted concepts, as shown in  LABEL:fig:intro_example. These shortages deviate text-based methods away from the following goals:
@@
A widely adopted erasing method, ESD (Gandikota et al., 2023), erases concepts by reducing $P(c|x)$, where they merely use a word to describe the concept $c$. However, one can easily design tricky yet benign prompts to generate NSFW content, failing to meet (G1). Another state-of-the-art method AdvUnlearn (Zhang et al., 2024b) successfully prevents the generation of undesirable content but is misaligned with benign text prompts (G2). This stems from the usage of a perturbed text description for the target concept, introducing implicit noise that may affect unrelated concepts. All these text-based erasing methods either fail to avoid unwanted generation given inductive prompts, or compromise generation quality given benign prompts, making it difficult to meet both goals simultaneously.
Figure 1: Performance overview of Co-Erasing and other methods when erasing nudity. A higher ASR indicates a lower erasing efficacy, meaning the target concept is not erased completely. A higher FID or a lower CLIP score indicates a lower general usability, meaning the benign generation is degraded. Competitive methods include FMN (Zhang et al., 2024a), ESD (Gandikota et al., 2023), SPM (Lyu et al., 2024), UCE (Gandikota et al., 2024), SalUn (Fan et al., 2024), SH (Wu & Harandi, 2024), ED (Jing et al., 2024) and AdvUnlearn (Zhang et al., 2024b).
@@ 2 Related Work
Concept Erasing. Current methods mainly follow two branches. One branch (Gandikota et al., 2023) (Huang et al., 2023) (Changhoon et al., 2024) (Chao et al., 2024) (Alvin & Harold, 2023) fine-tunes the entire model to adjust the output distribution away from the erased concept. UCE (Gandikota et al., 2024) erases in an image-editing way and manages to shift the output from the target concept to another modified concept. FMN (Zhang et al., 2024a) removes the concept-related knowledge by suppressing the attention paid to the target concept. SLD (Schramowski et al., 2023) introduces safety guidance during the iteration, similar to the classifier-free guidance (Ho, 2022). However, they often fail against learnable prompts, where semantically benign prompts can trick the model into generating undesirable content.
@@ 4.2.1 Integrating Image with Text Prompts
To introduce images into erasing, we require the image features to align with the text features in terms of dimension in order to collaboratively guide the diffusion process. Furthermore, we expect the image prompts to provide concept-related knowledge instead of low-level visual features such as structure and texture. Therefore, inspired by  (Ye et al., 2023), we propose to employ a decoupled cross-attention, where the text embedding $c_{\text{text}}\in\mathbb{R}^{b\times 77\times 768}$ and image embedding $c_{\text{image}}\in\mathbb{R}^{b\times 4\times 768}$ are separately fed into the layers, where $b$ is the batch size:
@@ 5.1 Experiment Setups
Tasks. Following the benchmark (Zhang et al., 2024b), we conduct the erasing in: (1) nudity, which aims to prevent the model from generating nudity-related content. (2) style, and we choose to erase the painting style of Van Gogh following most previous works. (3) objects. We erase objects including parachute, church, tench. We also validate the effectiveness of our method in erasing portraits and multi-concepts, with further details and results deferred to the appendix due to space limitations.
@@
Evaluation Setups. We evaluate with (1) pre-ASR (2) ASR (Zhang et al., 2024c) (3) P4D (Zhi-Yi et al., 2024) (4) CCE (Minh et al., 2023) (5) FID (Heusel et al., 2017) and (6) CLIP (Hessel et al., 2021). The former four measure the model’s ability to defend against manually designed and learnable prompts that aim to induce unwanted content, with a lower value indicating a more complete erasing. FID and CLIP score measure the model’s ability to generate normal content given benign prompts. A lower FID and a higher CLIP score indicate better general usability. Details about the metrics are deferred to  Section A.3.
Competitors. We include 9 competitors including (1) ESD (Gandikota et al., 2023), (2) FMN (Zhang et al., 2024a), (3) AC (Kumari et al., 2023), (4) UCE (Gandikota et al., 2024), (5) SPM (Lyu et al., 2024), (6) SH (Wu & Harandi, 2024), (7) ED (Jing et al., 2024), (8) SalUn (Fan et al., 2024), (9) AdvUnlearn (Zhang et al., 2024b). Note that some methods are not designed for all three tasks, and are evaluated on tasks they were originally employed.
@@
Figure 9: Examples of different methods erasing nudity against adversarial prompts. Prompts listed below images act as the base prompts before UDA (Zhang et al., 2024c).
@@ 5.2 Overall Performance
Co-Erasing can improve erasing efficacy. As shown in Figure 8, one can easily trick early methods including SPM, UCE, FMN, and ESD to generate unwanted concepts, as illustrated by their poor performance on efficacy. Specifically, compared to our baseline ESD, integrating images significantly improves efficacy when erasing nudity, which justifies the effectiveness of introducing image modality for erasing. In the first example in  Figure 9, it describes a woman in the oil painting style. ESD and FMN failed to prevent NSFW content from generation, while our generation can successfully avoid NSFW content.
@@
Co-Erasing can be transferred to other backbones. As our method is not restricted by the backbones, we apply Co-Erasing to other erasing frameworks such as SLD (Schramowski et al., 2023). Generally, when prompted by I2P dataset, the number of generations classified as nudity drops from $125$ to $22$. We defer implementation details and specific results to  Section B.5.
@@ 5.3 Quantative Analysis
Image modality is critical. As shown in the first two rows in  Table 1, ASR drops significantly from 76.05% to 32.60%, which strongly justifies the motivation and rationality of our method. However, the quality of the generation is degraded, with FID rising over $20$. This further requires us to keep unrelated concepts intact to preserve general usability.
@@ A.5 Evaluation Details
Most prompts for evaluation are from  (Zhang et al., 2024b), which are generated by GPT-4 and have been validated to be inductive. For experiments that are not conducted by  (Zhang et al., 2024b), we use GPT-4 to generate prompts, with some examples shown in  Table 5 and  Table 6.
@@
When evaluating style, we use an artistic style classification provided by  (Zhang et al., 2024c).
@@ B.5 Additional Backbones
To examine the effectiveness, we follow  Section B.4 to conduct a fine-grained analysis when erasing nudity by combining our method with SLD, as shown in  Table 18.
Table 18: Results of additional backbone when erasing nudity.
Exposed body part SLD SLD $+$ Co-Erasing
FEMALE_BREAST_EXPOSED 48 12
FEMALE_GENITALIA_EXPOSED 6 0
MALE_BREAST_EXPOSED 43 8
MALE_GENITALIA_EXPOSED 14 0
BUTTOCKS_EXPOSED 14 2
ANUS_EXPOSED 0 0

### v2

@@ 1 Introduction
Recent years have witnessed significant progress in the field of generation (Kingma & Welling, 2013; Goodfellow et al., 2014; Mirza & Osindero, 2014; Sohn et al., 2015), especially the diffusion model (Ho et al., 2020; Song et al., 2021). For example, the popular stable diffusion model  [^1] https://github.com/CompVis/stable-diffusion, trained on the LAION-5B dataset (Schuhmann et al., 2022), can produce high-quality images aligned with a given text prompt and have overwhelmed the area of image reconstruction (Takagi & Nishimoto, 2023; Wu et al., 2024; Li et al., 2023; Yang et al., 2021), edition (Meng et al., 2022; Yang et al., 2023a, b; Kawar et al., 2023; Bao et al., 2019) and generation (Rombach et al., 2022; Ho et al., 2022; Zhou et al., 2023; Bao et al., 2025; Zhang et al., 2024b). Despite its impressive generation ability, the diffusion model also raises safety concerns, one of which is the potential to generate unwanted content, such as NSFW (Not Safe For Work) images. Taking the SD v1.4 in  LABEL:fig:intro_example as an example, the designed prompt is semantically benign but induces explicit output.
To address this issue, a wide range of studies (Gandikota et al., 2023; Zhang et al., 2024a; Wu & Harandi, 2024; Schramowski et al., 2023; Kumari et al., 2023; Gandikota et al., 2024; Lyu et al., 2024; Zhang et al., 2024c; Fan et al., 2024; Ye et al., 2025; Huang et al., 2023; Bui et al., 2024; Changhoon et al., 2024; Chao et al., 2024; Lu et al., 2024; Alvin & Harold, 2023; Ye & Lu, 2023; Gao et al., 2024) have explored concept erasing techniques, which rely solely on text as guidance. However, as shown in this paper, their performances are limited by the inherent gap between the text and image modalities. Concepts are often complex and deeply entangled, making it challenging to fully separate with only text descriptions. Therefore, merely using text easily results in an incomplete erasing and affects other untargeted concepts, as shown in  LABEL:fig:intro_example. These shortages deviate text-based methods away from the following goals:
@@
A widely adopted erasing method, ESD (Gandikota et al., 2023), erases concepts by reducing $P(c|x)$, where they merely use a word to describe the concept $c$. However, one can easily design tricky yet benign prompts to generate NSFW content, failing to meet (G1). Another state-of-the-art method, AdvUnlearn (Zhang et al., 2024c), successfully prevents the generation of undesirable content but is misaligned with benign text prompts (G2). This stems from the usage of a perturbed text description for the target concept, introducing implicit noise that may affect unrelated concepts. All these text-based erasing methods either fail to avoid unwanted generation given inductive prompts, or compromise generation quality given benign prompts, making it difficult to meet both goals simultaneously.
Figure 1: Performance overview of Co-Erasing and other methods when erasing nudity. A higher ASR indicates a lower erasing efficacy, meaning the target concept is not erased completely. A higher FID or a lower CLIP score indicates a lower general usability, meaning the benign generation is degraded. Competitive methods include FMN (Zhang et al., 2024a), ESD (Gandikota et al., 2023), SPM (Lyu et al., 2024), UCE (Gandikota et al., 2024), SalUn (Fan et al., 2024), SH (Wu & Harandi, 2024), ED (Jing et al., 2024) and AdvUnlearn (Zhang et al., 2024c).
@@ 2 Related Work
Concept Erasing. Current methods mainly follow two branches. One branch (Gandikota et al., 2023) (Huang et al., 2023) (Changhoon et al., 2024) (Chao et al., 2024) (Alvin & Harold, 2023) (Ye & Lu, 2024; Bao et al., 2022) fine-tunes the entire model to adjust the output distribution away from the erased concept. UCE (Gandikota et al., 2024) erases in an image-editing way and manages to shift the output from the target concept to another modified concept. FMN (Zhang et al., 2024a) removes the concept-related knowledge by suppressing the attention paid to the target concept. SLD (Schramowski et al., 2023) introduces safety guidance during the iteration, similar to the classifier-free guidance (Ho, 2022). However, they often fail against learnable prompts, where semantically benign prompts can trick the model into generating undesirable content.
@@ 4.2.1 Integrating Image with Text Prompts
To introduce images into erasing, we require the image features to align with the text features in terms of dimension in order to collaboratively guide the diffusion process. Furthermore, we expect the image prompts to provide concept-related knowledge instead of low-level visual features such as structure and texture. Therefore, inspired by  (Ye et al., 2023; Li et al., 2024; Han et al., 2024), we propose to employ a decoupled cross-attention, where the text embedding $c_{\text{text}}\in\mathbb{R}^{b\times 77\times 768}$ and image embedding $c_{\text{image}}\in\mathbb{R}^{b\times 4\times 768}$ are separately fed into the layers, where $b$ is the batch size:
@@ 5.1 Experiment Setups
Tasks. Following the benchmark (Zhang et al., 2024c), we conduct the erasing in: (1) nudity, which aims to prevent the model from generating nudity-related content. (2) style, and we choose to erase the painting style of Van Gogh following most previous works. (3) objects. We erase objects including parachute, church, tench. We also validate the effectiveness of our method in erasing portraits and multi-concepts, with further details and results deferred to the appendix due to space limitations.
@@
Figure 9: Examples of different methods erasing nudity against adversarial prompts. Prompts listed below images act as the base prompts before UDA (Zhang et al., 2024d).
Evaluation Setups. We evaluate with (1) pre-ASR (2) ASR (Zhang et al., 2024d) (3) P4D (Zhi-Yi et al., 2024) (4) CCE (Minh et al., 2023) (5) FID (Heusel et al., 2017) and (6) CLIP (Hessel et al., 2021). The former four measure the model’s ability to defend against manually designed and learnable prompts that aim to induce unwanted content, with a lower value indicating a more complete erasing. FID and CLIP score measure the model’s ability to generate normal content given benign prompts. A lower FID and a higher CLIP score indicate better general usability. Details about the metrics are deferred to  Section A.3. We also consider a red-team method, RAB (Ring-A-Bell) (Tsai et al., 2023), to validate the robustness of Co-Erasing.
Competitors. We include 9 competitors including (1) ESD (Gandikota et al., 2023), (2) FMN (Zhang et al., 2024a), (3) AC (Kumari et al., 2023), (4) UCE (Gandikota et al., 2024), (5) SPM (Lyu et al., 2024), (6) SH (Wu & Harandi, 2024), (7) ED (Jing et al., 2024), (8) SalUn (Fan et al., 2024), (9) AdvUnlearn (Zhang et al., 2024c). Note that some methods are not designed for all three tasks, and are evaluated on tasks they were originally employed.
Table 1: Multi-concepts Erasure: Performance comparison when merged with MACE. $\text{Acc}_{e}$ and $\text{Acc}_{s}$ indicate the classification accuracy of the target and untargeted concepts, respectively. A higher $\text{H}_{c}=\frac{2}{(1-\text{Acc}_{e})^{-1}+\text{Acc}_{s}^{-1}}$ indicates a better performance. $\text{CLIP}_{e}$ and $\text{CLIP}_{s}$ indicate the classification accuracy of the target and untargeted concepts, respectively. A higher $\text{H}_{a}=\text{CLIP}_{s}-\text{CLIP}_{e}$ indicates a better performance.
Targets Objects
Dog Cat Bird Fish Horse
Method $\text{Acc}_{e}$ $\text{Acc}_{s}$ $\text{H}_{c}\uparrow$ $\text{Acc}_{e}$ $\text{Acc}_{s}$ $\text{H}_{c}\uparrow$ $\text{Acc}_{e}$ $\text{Acc}_{s}$ $\text{H}_{c}\uparrow$ $\text{Acc}_{e}$ $\text{Acc}_{s}$ $\text{H}_{c}\uparrow$ $\text{Acc}_{e}$ $\text{Acc}_{s}$ $\text{H}_{c}\uparrow$
SD 98.250 99.000 0.034 98.500 98.800 0.030 96.720 99.500 0.064 97.120 99.200 0.056 98.450 99.250 0.031
MACE 6.370 86.885 0.901 10.060 94.328 0.921 8.000 97.955 0.949 0.870 94.628 0.968 5.880 94.355 0.942
MACE+Co-Erasing 4.620 89.408 0.923 8.730 94.308 0.928 7.800 98.173 0.951 0.790 94.033 0.966 5.550 94.110 0.943
Targets Objects Celebrities Style
5 Objects 1 Celebritiy 10 Celebrities 100 Celebrities 100 Artists
Method $\text{Acc}_{e}$ $\text{Acc}_{s}$ $\text{H}_{c}\uparrow$ $\text{Acc}_{e}$ $\text{Acc}_{s}$ $\text{H}_{c}\uparrow$ $\text{Acc}_{e}$ $\text{Acc}_{s}$ $\text{H}_{c}\uparrow$ $\text{Acc}_{e}$ $\text{Acc}_{s}$ $\text{H}_{c}\uparrow$ $\text{CLIP}_{e}$ $\text{CLIP}_{s}$ $\text{H}_{a}\uparrow$
SD 99.000 98.650 0.020 92.200 94.420 0.144 96.350 94.420 0.070 95.660 94.420 0.083 31.220 31.220 -
MACE 7.800 12.570 0.221 0.500 93.550 0.964 2.910 92.510 0.947 4.520 85.110 0.900 22.140 27.790 5.650
MACE+Co-Erasing 6.880 13.260 0.232 0.000 92.960 0.964 2.790 92.370 0.947 3.140 83.880 0.899 20.650 27.710 7.060
@@ 5.2 Overall Performance
Co-Erasing can improve erasing efficacy. As shown in Figure 8, one can easily trick early methods including SPM, UCE, FMN, and ESD to generate unwanted concepts, as illustrated by their poor performance on efficacy. Specifically, compared to our baseline ESD, integrating images significantly improves efficacy when erasing nudity, which justifies the effectiveness of introducing image modality for erasing. In the first example in  Figure 9, it describes a woman in the oil painting style. ESD and FMN failed to prevent NSFW content from generation, while our generation can successfully avoid NSFW content. Furthermore, we merge our Co-Erasing with a widely used multi-concept erase framework MACE (Lu et al., 2024) as shown in  Table 1, which again validates the efficacy and generalization.
@@
Co-Erasing can be transferred to other backbones. As our method is not restricted by the backbones, we apply Co-Erasing to other erasing frameworks such as SLD (Schramowski et al., 2023) and MACE (Lu et al., 2024). When evaluated by RAB (Tsai et al., 2023), all backbones can effectively improve erasing efficacy, as shown in  Table 2. Specifically, when prompted by the I2P dataset, the number of SLD generations classified as nudity drops from $125$ to $22$. We defer implementation details and specific results to  Section B.5.
Table 2: Performance evaluation against Ring-A-Bell.
Method RAB_K16 $\downarrow$ RAB_K38 $\downarrow$ RAB_K77 $\downarrow$ FID $\downarrow$ CLIP $\uparrow$
SD 0.74 0.78 0.68 14.77 0.312
ESD 0.47 0.52 0.42 18.18 0.302
ESD+Co-Erasing 0.18 0.22 0.22 18.77 0.302
MACE 0.04 0.02 0.00 17.13 0.277
MACE+Co-Erasing 0.00 0.00 0.00 17.12 0.277
@@ 5.3 Quantative Analysis
Table 3: Validation on text, image, and text-guided refinement.
@@
Image modality is critical. As shown in the first two rows in  Table 3, ASR drops significantly from 76.05% to 32.60%, which strongly justifies the motivation and rationality of our method. However, the quality of the generation is degraded, with FID rising over $20$. This further requires us to keep unrelated concepts intact to preserve general usability. To further understand how image modality helps, we use visualizations to show that images can fill in semantic gaps left by sparse or ambiguous text. We use LLM to generate an expression set (including words and phrases) related to the concept of nudity. With a technique similar to Textual-Inversion, we use model-generated images to retrieve the top-5 related expressions from the set, as shown in  Figure 11. The retrieved phrases from images often reflect a broader and more nuanced conceptual space than the original seed terms (e.g., “nudity”, “sexy”). This suggests that image embeddings encode richer semantic representations than their originating textual prompts.
Figure 11: Retrieved Top-5 phrases by generated images. Note that all images are refined by text-guidance “nudity” before retrieving. Results show that images can reflect a broader conceptual space to recover latent semantics.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by the National Key R&D Program of China under Grant 2018AAA0102000, in part by National Natural Science Foundation of China: 62236008, 62441232, 62441619, U21B2038, U23B2051, 62122075, 62206264 and 92370102, in part by the Open Topics from the Lion Rock Labs of Cyberspace Security (under the project #LRL24009), in part by the Youth Innovation Promotion Association CAS, in part by the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB0680201, in part by the Fundamental Research Funds for the Central Universities, and in part by the Postdoctoral Fellowship Program of CPSF under Grant GZB20240729.
@@ A.5 Evaluation Details
Most prompts for evaluation are from  (Zhang et al., 2024c), which are generated by GPT-4 and have been validated to be inductive. For experiments that are not conducted by  (Zhang et al., 2024c), we use GPT-4 to generate prompts, with some examples shown in  Table 7 and  Table 8.
@@
When evaluating style, we use an artistic style classification provided by  (Zhang et al., 2024d).
@@ B.5 Additional Backbones
To examine the effectiveness, we follow  Section B.4 to conduct a fine-grained analysis when erasing nudity by combining our method with SLD, ESD, and MACE, as shown in  Table 20.
Table 20: Assessment of Explicit Content Removal: F: Female. M: Male. Prompts are from I2P benchmark, and SD v1.4 serves as the baseline for all methods.
Method NudeNet Detection COCO-10K
Breasts(F) Breasts(M) Genitalia(F) Genitalia(M) Buttocks Anus Total $\downarrow$ FID $\downarrow$ CLIP $\uparrow$
SLD 48 43 6 14 14 0 125 20.17 0.298
SLD+Co-Erasing 12 8 0 0 2 0 22 21.15 0.297
ESD 27 8 3 2 2 1 43 18.18 0.302
ESD+Co-Erasing 2 1 0 0 1 0 4 18.77 0.302
MACE 16 9 2 7 2 0 36 17.13 0.277
MACE+Co-Erasing 4 2 1 1 1 0 9 17.12 0.277